Full-time students at a state school in New England earned between 10 and 16 credits in Spring 2012. The table below gives the distribution of the number of credits earned. # of credits 10 11 12 13 14 15 16 % of students 1 3 6 28 33 21 8 Find the approximate probability that the total number of credits earned by a random sample of 484 students from that school in that semester was more than 6750.
Answer #1

for a single student:

x P(x) xP(x) x2P(x)
10 0.01 0.100 1.000
11 0.03 0.330 3.630
12 0.06 0.720 8.640
13 0.28 3.640 47.320
14 0.33 4.620 64.680
15 0.21 3.150 47.250
16 0.08 1.280 20.480
total 13.840 193.000
E(x) =μ= ΣxP(x) = 13.8400
E(x2) = Σx2P(x) = 193.0000
Var(x)=σ2 = E(x2)-(E(x))2= 1.454
std deviation=         σ= √σ2 = 1.2060

therefore for 484 students expected mean =484*13.84= 6698.56

and standard deviation =1.2060*sqrt(484)=26.532

therefore from normal approximation;approximate probability that the total number of credits earned by a random sample of 484 students from that school in that semester was more than 6750

=P(X>6750)=P(Z>(6750-6698.56)/26.532)=P(Z>1.94)=1-P(Z<1.94)=1-0.9738 =0.0262